# Pretraining Frame Preservation in Autoregressive Video Memory Compression

### v3

@@ Abstract
We present a neural network structure to compress long videos into short contexts, with an explicit pretraining objective to preserve the high-frequency details of single frames at arbitrary temporal positions. The baseline model can compress a 20-second video into a context at about 5k length, where random frames can be retrieved at perceptually preserved appearances. Such pretrained models can be directly fine-tuned as memory encoders for autoregressive video models, enabling long history memory with low context cost and relatively low fidelity loss. We evaluate the framework with ablative settings and discuss the trade-offs of possible neural architecture designs.
Figure 1: This video was generated autoregressively second-by-second with full history context (without cutting out any history frames). The 20+ second history was compressed into a $\sim$5k context length and processed by an RTX 4070 12GB.
@@ 1 Introduction
Storytelling capability, narrative coherence, and context quality are increasingly important in the video generation community. The latest commercial models like Sora2 [46], Veo3.1 [20], and Kling2.5 [37] enable storyboard creation-based, scene planning, and dynamic camera shots (e.g., in the style of TikTok or YouTube Shorts). Recent academic models have also focused on long video streaming [72, 27], scene planning [22, 26, 83, 86, 87], and video context consistency [28, 31, 42]. Among these approaches, autoregressive models are a key paradigm for video storytelling, supporting native video continuation and storyboard streaming. Autoregressive models view video history as context, and they face unique challenges when handling long-form content like in-the-wild videos and movies.
@@
The trade-off between context quality and length is a typical and critical problem in video context modeling. A naive sliding window, which cuts off all distant-enough frames, maintains a fixed context length but loses long-range history. Various compression approaches have been extensively studied; e.g., highly compressed VAEs like LTXV [23] and DC-AE [10] show that it is possible to compress pixel-space videos into more compact contexts, and hybrid approaches [1, 32, 88] like FramePack [78] show that compression can happen at multiple levels, albeit at the cost of high-frequency image details. Another strategy is to preserve the context length but reduce computation (e.g., using sparse [64, 77, 80, 65] or linear [5, 68, 58, 11, 74, 33] attention), though the linear layer cost remains an overhead for training and bidirectional inference. More specialized methods like token merging [4, 3] also demonstrate that a higher merging rate leads to greater detail loss. These research efforts implicitly converge on the observation that context quality and length are governed by a fundamental trade-off where the optimal balance has not yet been reached.
This observation leads to a wide range of open questions. Is there an explicit correlation between context length and quality, and how can we model it? Should context quality be measured by object, scene, or facial consistency, or potentially, by full-frame reconstruction quality? To achieve a perceptually acceptable quality from a human perspective, what is the maximum level of compression, and where is the sweet spot? How can such a balance be found? Do different tasks and objectives pose varied requirements for this trade-off? Do spatial and temporal compression have different levels of importance in various tasks? Is there a marginal effect of context length that makes a sufficiently long context actually redundant?
To study these questions, this paper targets an explicit objective: a learning task to directly compress video history frames into a minimized context length while maximizing frame retrieval quality. This setting allows us to explicitly study the correlation between context length and reconstruction quality using both quantitative metrics and human perceptual evaluations. We can thereby independently study the quality, influence, and cost of the compression model before moving on to the actual autoregressive video generation. This white-box feedback eases cost management and contributes to the explainability of separate encoding components in autoregressive video models.
To this end, one challenge is that learning to compress and transform a long video context causes computational bottlenecks. We observe that a strong indicator of a video compression mechanism’s ability to preserve contextual details is its capacity for high-quality frame reconstruction at arbitrary temporal locations. This leads to a randomized approach: we first pretrain a memory encoder to reconstruct frames at random time positions, and then finetune an autoregressive video model using the compression model as a history memory encoder. Pretraining with millions of videos, we show that this method reduces the video model training cost by a large margin and facilitates a stronger neural representation for detailed frame retrieval.
@@
Experiments show that the pretrained memory encoder can, by itself, reconstruct history frames at high quality (measured by PSNR/SSIM [60]). When connected to video diffusion models for autoregressive finetuning, the encoder provides long-range consistency. This framework achieves a balanced and optimized point in the context length-quality trade-off, leading to a practical autoregressive video model with a relatively short context length and perceptually high consistency based on its long-range history.
@@
In summary, we (1) discuss the trade-off between context length and quality in autoregressive video generation; (2) propose to pretrain an explicit compression task for frame retrieval for improved autoregressive video memory; (3) establish an autoregressive system to fine-tune the pretrained compression model as a history memory encoder for DiTs; (4) conduct extensive experiments, including quantitative, qualitative, and ablation studies, to analyze the influence of different compression designs; and (5) provide pretrained models for practical video generation and editing applications.
@@ 2 Related Work
Context learning and compression. Long-context persistence is crucial as we extend video generation beyond a few seconds. One option is retrieval, which grounds generation in a persistent state. WorldMem [67] and Context-as-Memory [73] augment AR video world models with FoV-based history retrieval, while VMem [39] indexes past views by surfels to retrieve the most relevant viewpoints. Memory Forcing [25] pairs geometry-indexed spatial memory with tailored training regimes to balance exploration and revisits. Beyond fixed retrieval rules, Mixture-of-Contexts [6] learns a dynamic sparse attention context router so that tokens attend only to the most salient chunks. Similarly, MoGA [30] and Holocine [45] also propose token-level sparse attending policies. Pack-and-Force [63] instead proposes a learnable context semantic retriever. Orthogonal to retrieval, compression of history turns an unbounded context into a compact state. FramePack [78] compresses prior frames into a fixed‑size latent “packed” context. Captain Cinema [66] uses a similar compression for keyframes. StateSpaceDiffuser [50] and Po et al. [47] swap quadratic attention for recurrent states to maintain long‑term memory. TTTVideo [13] and LaCT [81] use light MLP test-time training layers as a learned context representation.
@@
Besides, this approach presents distinctions with VAEs: VAEs have bottlenecks with 4, 16, or 64 channels, whereas this method directly works at DiT inner channels like 3072 or 5120 with a different feature scope; VAEs have training overheads to process full contexts, while this method works with lightweight partial overheads (retrieval objective).
@@ 3 Method
We pretrain a video compression model with the objective to retrieve frames at arbitrary time positions, and simultaneously preserve frame details (Section 3.1). The pretrained compression model is then finetuned as a memory encoder for autoregressive video diffusion models (Section 3.2).
@@
from clean latents $\bm{X}_{0}$ where $t_{i}\in(0,1]$ is the diffusion timestep. The autoregressive model typically takes the history $\bm{H}\in\mathbb{R}^{T_{h}\times H\times W\times C}$ as conditioning. To learn the autoregressive generator $\bm{G}_{\theta}(\cdot)$, the learning objective is
$\vskip-3.0pt\mathbb{E}_{\bm{X}_{0},\bm{H},\bm{c},\bm{\epsilon},t_{i}\sim\mathcal{L}(0,1)}\bigg\|(\bm{\epsilon}-\bm{X}_{0})-\bm{G}_{\theta}\big(\bm{X}_{t_{i}},t_{i},\bm{c},\bm{H}\big)\bigg\|_{2}^{2},$ (2)
@@
3.1 Pretraining Memory Compression Model
Figure 2: Pretraining of memory compression models. The memory compression model has to compress long videos (e.g., 20 seconds) into short contexts (e.g., of length 5k). The objective of the pretraining is to retrieve frames with high-frequency details in arbitrary history time positions.
@@ 3.1 Pretraining Memory Compression Model
We observe that a well-suited indicator of a video compression mechanism’s ability to preserve contextual details is its capacity for high-quality frame retrieval at arbitrary time positions. For high compression rates, perfect retrieval becomes impractical, leading to a goal to maximize the retrieval quality of arbitrary frames.
We formulate this goal in a technical form. Considering that a compression process $\phi(\cdot)$ receives a history $\bm{H}$ of $T_{h}$ frames and compresses it into a context $\phi(\bm{H})$, given an arbitrary set of frame indices $\bm{\Omega}\subseteq\mathbb{N}$, and denoting a indexed retrieval process as $\phi^{-1}(\cdot,\bm{\Omega})$, we want to minimize the feature distance $\|\bm{H}_{\bm{\Omega}}-\phi^{-1}(\phi(\bm{H}),\bm{\Omega})\|_{2}^{2}$ for optimized reconstruction.
As shown in Fig. 2, the frame selection $\bm{\Omega}$ and retrieval $\phi^{-1}(\cdot,\bm{\Omega})$ can be established as an autoregressive video diffusion framework. We sample random videos as the history, and then pick random frames $\bm{\Omega}$: in particular, we randomly keep a subset of frames at indices $\bm{\Omega}$ unchanged and then mask all remaining frames. Various types of masks can be utilized like range scaling, shifting, etc. Herein, we use a noise-as-mask method: add latent noise levels from $\mathcal{L}(0.2,1)$. After that, we clone the picked clean frames as the diffusion target, so that the diffusion system will try to reconstruct the target frames at arbitrary positions. This process can also be written in line with Eq. (2) as
@@
$\vskip-3.0pt\mathbb{E}_{\bm{H},\bm{\Omega},\bm{c},\bm{\epsilon},t_{i}}\bigg\|(\bm{\epsilon}-\bm{H}_{\bm{\Omega}})-\bm{G}_{\theta}\big((\bm{H}_{\bm{\Omega}})_{t_{i}},t_{i},\bm{c},\phi(\bm{H})\big)\bigg\|_{2}^{2},$ (3)
where $\bm{G}_{\theta}(\cdot)$ can be a trainable video diffusion model like WAN or HunyuanVideo with LoRAs.
@@
When the output length of the compression model $\phi(\cdot)$ is short, e.g., about only 5k length for every 20 seconds, the retrieval task becomes highly difficult, making it well-suited for pretraining strong context compression models using millions of in-the-wild videos. We also point out the importance of the randomness in frame indices $\bm{\Omega}$ selection: if the frames are not randomly selected, e.g., only training with fixed starting or ending positions, the framework will learn a cheating solution, e.g., using all 5k context to encode a few ending frames while ignoring all other frames. Besides, note that directly training to retrieve all frames is impractical with minute-level data, even for highly efficient infrastructures.
@@
Figure 3: Architecture of memory compression model. We use 3D convolution, SiLU, and attention to establish a lightweight neural structure as the baseline compression model. Different alternative architectures (e.g., various channels, full transformer, etc.) are possible and will be discussed in ablation.
@@
Network architecture. As shown in Fig. 3, we present a lightweight baseline architecture for the compression model $\phi(\cdot)$. Instead of building the compressed representation from scratch, we reuse the DiT’s context representation: we first downsample a high-resolution high-fps video into a low-resolution low-fps one and process it with the VAEs and DiT’s patchifiers and first-layer projections. After that, we encode the high-resolution original video into a residual-enhancing vector and add it to the context vector. Note that the feature addition happens after the DiT’s first-layer projection, which means the encoder does not go through the VAE’s 16 channel bottleneck but directly outputs at inner channels (like 3072 for WAN-5B) for best fidelity. The baseline encoder uses 3D convolutions as starting layers and ends with attention layers. Besides, it is also possible to connect the DiT with cross-attentions from the last hidden states of the encoder for another refinement pass, but the quality gain is marginal and will be covered in ablations.
@@ 3.2 Finetuning Video Diffusion Model
Figure 4: Finetuning autoregressive video models. We illustrate the finetuning and inference of the final autoregressive video models. The pretraining of the memory compression model is finished before the finetuning.
With the pretrained memory compression model $\phi(\cdot)$, we can form autoregressive video generation systems by finetuning video diffusion models like WAN (with LoRAs) and the pretrained compression model as history memory encoders. This will lead to an autoregressive video model with long history windows (e.g., >> 20s), short history context length (e.g., about 5k), and explicitly optimized frame retrieval quality. The diffusion can be written in line with Eq (2) as
$\vskip-3.0pt\mathbb{E}_{\bm{X}_{0},\bm{H},\bm{c},\bm{\epsilon},t_{i}}\bigg\|(\bm{\epsilon}-\bm{X}_{0})-\bm{G}_{\theta}\big(\bm{X}_{t_{i}},t_{i},\bm{c},\phi(\bm{H})\big)\bigg\|_{2}^{2},$ (4)
where the process is as shown in Fig. 4-(a). The inference follows autoregressive generation by repeatedly concatenating the generated sections into the history as shown in Fig. 4-(b). Note that since the baseline encoder (Fig. 3) is almost fully convolutional, the compressed history can also be concatenated on-the-fly without recomputation.
Figure 5: Visual comparison on compression reconstruction. We present reconstruction results after the pretraining using different possible neural structures and various compression settings. * The “Large Patchifier” is technically equivalent to [78].
@@ 4.1 Experimental Details
Hyper-parameters. When constructing the encoders of the memory compression models, we use the denotation $H\times W\times T$ to mark the compression rates, e.g., $4\times 4\times 2$ means compressing the latent width 4 times, height 4 times, and temporal length 2 times. Note that this compression rate is on top of the latent space and can be multiplied by the latent rate and patchifier rate to compute the final compression rate against the pixel space. When building the 3D convolution layers, we follow the rule to always first reduce the temporal dimension and then the spatial dimension (using strides in convolutions). The hidden channels of the encoder follow the pattern $64\rightarrow 128\rightarrow 256\rightarrow 512$ and are then always kept at $512$ (e.g., in the final attention layer) before the last $1\times 1$ convolution, which projects the feature to the same dimensions (like 3072, 5120, etc.) as the DiT context.

### v5

# Pretraining Frame Preservation for Lightweight Autoregressive Video History Embedding

@@ Abstract
Autoregressive video generation relies on history context for content consistency and storytelling. As video histories grow longer, efficiently encoding them remains an open problem — particularly for personal users and local workflows where compute and memory budgets are limited. We present a lightweight history encoder that maps long video histories into short-length embeddings, pretrained with a frame query objective that learns to attend to content features at arbitrary temporal positions. The pretraining stage provides the encoder with dense history coverage on large-scale video data; the subsequent finetuning stage adapts the pretrained encoder under an autoregressive video generation objective to establish content-level consistency. In this way, the lightweight embeddings achieve comparable performance to heavier alternatives. We evaluate the framework with ablative settings and discuss the architecture designs.
Figure 1: This video was generated autoregressively second-by-second with full history context (without cutting out any history frames). The full history is encoded into a lightweight embedding and processed by an RTX 4070 12GB.
@@ 1 Introduction
Storytelling capability, narrative coherence, and context quality are increasingly important in the video generation community. The latest commercial models like Sora2 [46], Veo3.1 [20], and Kling2.5 [37] enable storyboard creation-based, scene planning, and dynamic camera shots (e.g., in the style of TikTok or YouTube Shorts). Recent academic models have also focused on long video streaming [71, 27], scene planning [22, 26, 82, 85, 86], and video context consistency [28, 31, 42]. Among these approaches, autoregressive models are a key paradigm for video storytelling, supporting native video continuation and storyboard streaming. Autoregressive models view video history as context, and they face unique challenges when handling long-form content. These challenges are amplified for personal users, offline pipelines, and local finetuning, where encoding longer video histories must fit within limited compute and memory.
@@
The trade-off between context quality and length is a typical and critical problem in video context modeling. A naive sliding window, which cuts off all distant-enough frames, maintains a fixed context length but loses long-range history. Various compression approaches have been extensively studied; e.g., highly compressed VAEs like LTXV [23] and DC-AE [10] show that it is possible to compress pixel-space videos into more compact contexts, and hybrid approaches [1, 32, 87] like FramePack [77] show that compression can happen at multiple levels, albeit at the cost of high-frequency image details. Another strategy is to preserve the context length but reduce computation (e.g., using sparse [63, 76, 79, 64] or linear [5, 67, 58, 11, 73, 33] attention), though the linear layer cost remains an overhead for training and bidirectional inference. More specialized methods like token merging [4, 3] also demonstrate that a higher merging rate leads to greater detail loss. These efforts indicate that the balance among history coverage, visual fidelity, and memory cost remains an open problem, and that lightweight history encoding is broadly desired.
Several questions arise in this setting. Should context quality be measured by object, scene, or facial consistency? Do spatial and temporal encoding have different levels of importance? Is there a diminishing return of context length beyond a certain point? A practical approach should address two concerns: learning to represent video history densely and efficiently, and learning to generate content that is consistent with that history. Separating these concerns into a pretraining stage and a finetuning stage reduces the computational cost compared to joint training, and enables each stage to use different data and objectives.
To this end, this paper presents a history encoder that maps video history into a lightweight embedding, learned through a pretraining-then-finetuning approach. In the pretraining stage, the encoder is trained on large-scale video data with a frame query objective: given a long history, the encoder must enable the diffusion model to attend to frames at arbitrary temporal positions. Only a random subset of frames is supervised at each step, ensuring dense history coverage while keeping training cost low. In the finetuning stage, the pretrained encoder is integrated into an autoregressive video diffusion model and adapted jointly to establish content-level consistency.
One challenge in learning such a history encoder is aligning its output with the generation model’s inner feature space. Instead of building the encoded representation from scratch, we reuse the DiT’s own processing pipeline: the encoder outputs directly at the DiT’s inner hidden states (e.g., 3072 or 5120 channels), bypassing the VAE’s narrow bottleneck. This design allows the encoder to manipulate deep DiT features rather than the latent space. Pretraining on millions of videos, this approach provides a strong initialization for the subsequent finetuning stage, reducing training cost and improving the quality of the learned history embedding.
@@
Experiments show that the pretrained encoder can, by itself, query history frames with high feature quality. When integrated into a video diffusion model for autoregressive finetuning, the encoder provides long-range content consistency. This framework leads to a practical autoregressive video model with a lightweight history context and perceptually high consistency, as validated by both quantitative evaluations and human studies.
@@
In summary, we (1) discuss the trade-off between context length and quality in autoregressive video generation, and the need for lightweight history encoding in personal and local workflows; (2) propose to pretrain a history encoder with a frame query objective, providing dense coverage and reducing subsequent finetuning cost; (3) integrate the pretrained encoder into an autoregressive video diffusion system and finetune for content-level consistency; (4) conduct extensive experiments, including quantitative, qualitative, and ablation studies, to analyze the influence of different architecture designs; and (5) provide pretrained models for practical video generation applications.
@@ 2 Related Work
Context learning and history encoding. Long-context persistence is crucial as we extend video generation beyond a few seconds. One option is retrieval, which grounds generation in a persistent state. WorldMem [66] and Context-as-Memory [72] augment AR video world models with FoV-based history retrieval, while VMem [39] indexes past views by surfels to retrieve the most relevant viewpoints. Memory Forcing [25] pairs geometry-indexed spatial memory with tailored training regimes to balance exploration and revisits. Beyond fixed retrieval rules, Mixture-of-Contexts [6] learns a dynamic sparse attention context router so that tokens attend only to the most salient chunks. Similarly, MoGA [30] and Holocine [45] also propose token-level sparse attending policies. Pack-and-Force [62] instead proposes a learnable context semantic retriever. Orthogonal to retrieval, compression of history turns an unbounded context into a compact state. FramePack [77] compresses prior frames into a fixed‑size latent “packed” context. Captain Cinema [65] uses a similar compression for keyframes. StateSpaceDiffuser [50] and Po et al. [47] swap quadratic attention for recurrent states to maintain long‑term memory. TTTVideo [13] and LaCT [80] use light MLP test-time training layers as a learned context representation.
@@
Besides, this approach presents distinctions with VAEs: VAEs have bottlenecks with 4, 16, or 64 channels, whereas this method directly works at DiT inner channels like 3072 or 5120 with a different feature scope; VAEs have training overheads to process full contexts, while this method works with lightweight partial overheads (frame query objective).
@@ 3 Method
We pretrain a history encoder with the objective to query content features at arbitrary time positions (Section 3.1). The pretrained encoder is then finetuned within autoregressive video diffusion models (Section 3.2).
@@
from clean latents $\bm{X}_{0}$ where $t_{i}\in(0,1]$ is the diffusion timestep. The autoregressive model typically takes the history $\bm{H}\in\mathbb{R}^{T_{h}\times H\times W\times C}$ as conditioning. The generator $\bm{G}_{\theta}(\cdot)$ is trained by finding
$\begin{split}\arg\min_{\theta}&\mathbb{E}_{\bm{X}_{0},\bm{H},\bm{c},\bm{\epsilon},t_{i}\sim\mathcal{L}(0,1)}\\ &\bigg\|(\bm{\epsilon}-\bm{X}_{0})-\bm{G}_{\theta}\big(\bm{X}_{t_{i}},t_{i},\bm{c},\bm{H}\big)\bigg\|_{2}^{2},\end{split}$ (2)
@@
3.1 Pretraining History Encoder
Figure 2: Pretraining the history encoder. The encoder maps long video histories into lightweight embeddings. The objective of the pretraining is to query content features at arbitrary history time positions.
@@ 3.1 Pretraining History Encoder
We observe that a useful indicator of the history encoder’s ability to preserve content features is its capacity to query frames at arbitrary temporal positions. Given a video history $\bm{H}$, we randomly sample a set of frame indices $\bm{\Omega}$. Frames at the selected indices are kept clean, while all remaining frames are masked with noise sampled from $\mathcal{L}(0.2,1)$, resulting in a masked history $\texttt{mask}(\bm{H},\bm{\Omega})$. As shown in Fig. 2, the DiT learns to query the selected frames via a conditional mapping
$\bm{G}_{\theta}:\{\phi(\texttt{mask}(\bm{H},\bm{\Omega})),\bm{c}\}\mapsto\bm{H}_{[\bm{\Omega}]},$ (3)
trained by finding
@@
$\begin{split}&\arg\min_{\theta,\phi}\mathbb{E}_{\bm{H},\bm{\Omega},\bm{c},\bm{\epsilon},t_{i}}\\ &\bigg\|(\bm{\epsilon}-\bm{H}_{[\bm{\Omega}]})-\bm{G}_{\theta}\big((\bm{H}_{[\bm{\Omega}]})_{t_{i}},t_{i},\bm{c},\phi(\texttt{mask}(\bm{H},\bm{\Omega}))\big)\bigg\|_{2}^{2},\end{split}$ (4)
where $\bm{G}_{\theta}(\cdot)$ is the DiT with low-rank adaptation (LoRA).
@@
When the output of the encoder $\phi(\cdot)$ is short relative to the original history, the frame query task becomes highly difficult, making it well-suited for pretraining history encoders on millions of in-the-wild videos. The randomness of $\bm{\Omega}$ is important: if only frames at fixed positions (e.g., the last few frames) are used as targets, the encoder can learn a cheating shortcut by dedicating all capacity to those positions while ignoring the rest of the history. Random selection forces the embedding to maintain dense coverage across the entire temporal extent. Besides, directly training to query all frames is impractical with minute-level data, even for highly efficient infrastructures.
@@
Figure 3: Architecture of the history encoder. We use 3D convolution, SiLU, and attention to establish a lightweight neural structure as the baseline encoder. Different alternative architectures (e.g., various channels, full transformer, etc.) are possible and will be discussed in ablation.
@@
Network architecture. As shown in Fig. 3, we present a lightweight baseline architecture for the history encoder $\phi(\cdot)$. Instead of building the compressed representation from scratch, we reuse the DiT’s context representation: we first downsample a high-resolution high-fps video into a low-resolution low-fps one and process it with the VAEs and DiT’s patchifiers and first-layer projections. After that, we encode the high-resolution original video into a residual-enhancing vector and add it to the context vector. Note that the feature addition happens after the DiT’s first-layer projection, which means the encoder does not go through the VAE’s 16 channel bottleneck but directly outputs at inner channels (like 3072 for WAN-5B) for best fidelity. The baseline encoder uses 3D convolutions as starting layers and ends with attention layers. Besides, it is also possible to connect the DiT with cross-attentions from the last hidden states of the encoder for another refinement pass, but the quality gain is marginal and will be covered in ablations.
@@ 3.2 Finetuning Video Diffusion Model
Figure 4: Finetuning autoregressive video models. We illustrate the finetuning and inference of the final autoregressive video models. The pretraining of the history encoder is finished before the finetuning.
With the pretrained history encoder $\phi(\cdot)$, we finetune the video diffusion model (with LoRA) together with the encoder for autoregressive generation. Unlike the pretraining stage, the history $\bm{H}$ is now complete and noise-free, and no frame indices $\bm{\Omega}$ are involved. The DiT learns to produce the next video segment $\bm{X}$ conditioned on the history embedding
$\bm{G}_{\theta}:\{\phi(\bm{H}),\bm{c}\}\mapsto\bm{X},$ (5)
by the diffusion objective
$\begin{split}\arg\min_{\theta,\phi}&\mathbb{E}_{\bm{X}_{0},\bm{H},\bm{c},\bm{\epsilon},t_{i}}\\ &\bigg\|(\bm{\epsilon}-\bm{X}_{0})-\bm{G}_{\theta}\big(\bm{X}_{t_{i}},t_{i},\bm{c},\phi(\bm{H})\big)\bigg\|_{2}^{2},\end{split}$ (6)
where newly initialized LoRA parameters are applied. As shown in Fig. 4-(a), the encoder $\phi(\cdot)$ continues to be updated during finetuning, allowing its features to adapt from frame-level query targets toward generation-oriented consistency. The inference follows autoregressive generation by repeatedly concatenating the generated sections into the history as shown in Fig. 4-(b). Since the encoder is almost fully convolutional, the encoded history can also be concatenated on-the-fly without recomputation.
Figure 5: Visual comparison on frame query results. We present results after the pretraining using different possible neural structures and various compression settings. * The “Large Patchifier” is technically equivalent to [77].
@@ 4.1 Experimental Details
[…]